OpenStudy (anonymous):

Solve for x. 8x − 5 = −21 ? show work

OpenStudy (anonymous):

add 5 to both sides

OpenStudy (anonymous):

then divide by 8

OpenStudy (saifoo.khan):

8x = -21 + 5 8x = -16 x = -16/2 x = 2